The half-life of a substance is 12 hours. Find k to complete the formula A = A sub 0 e^kt. After every computation, round to the nearest hundredth.
Thanks in advance for your help.

Answer by Cromlix(4381)

Hi there,
A = Ao e^kt
Starting wit 10 kg
Activity drops to half in 12 hours
5 = 10 e^k12
5/10 = e^12k
0.5 = e^12k
loge0.5 = 12k loge e (loge e = 1)
k = loge 0.5/12
k = -0.06
Hope this helps :-)